OpenStudy (anonymous):

solve. │x – 5│ = 3 i need help solving this

jimthompson5910 (jim_thompson5910):

Hint: |x-5| = 3 turns into x-5=3 or x-5 = -3

jimthompson5910 (jim_thompson5910):

solve each for x to get your two answers

OpenStudy (anonymous):

x = 8 and x = –2

jimthompson5910 (jim_thompson5910):

first one is correct second is not

OpenStudy (anonymous):

should it be a positive 2?

jimthompson5910 (jim_thompson5910):

yep, visually, if you draw out a number line 

jimthompson5910 (jim_thompson5910):

|x - 5| = 3 asks the question: which numbers x are 3 units away from the number 5? The answer is x = 2 or x = 8

jimthompson5910 (jim_thompson5910):